### v1

# Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages

Abstract
The current Large Language Models (LLMs) face significant challenges in improving performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning. From the perspective of language-bridge, we propose BridgeX-ICL, a simple yet effective method to improve zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages. Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs or not. We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly, to ensure full activation of these anchored neurons. Subsequently, we propose an HSIC-based metric to quantify LLMs-internal linguistic spectrum based on overlap neurons, which guides optimal bridge selection. The experiments conducted on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families (covering both high-low and moderate-low pairs) validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs.
1 Introduction
Although Large Language Models (LLMs) have demonstrated impressive multilingual capacities, there is still significant space for improving the performance on low-resource languages Huang et al. (2024); Nazi et al. (2025). To address this issue, especially avoiding costly post-training Muller et al. (2021); Yong et al. (2023), it is critical to fully investigate the multilingual understanding and transferring ability in LLMs.
Recent research has increasingly focused on data-efficient methods, particularly Cross-lingual In-Context Learning (X-ICL) Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024), which surprisingly works well on low-resource languages, primarily due to LLMs are in-context low-resource language learners Brown et al. (2020b); Cahyawijaya et al. (2024). For instance, in the Arabic-to-Hebrew Bilingual Lexicon Induction (BLI) task, the zero-shot baseline accuracy in LLaMA 3 is 47.0%. However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which significantly outperforms even the two-shot X-ICL. This observation motivates us to further explore: How can we improve cross-lingual capabilities of LLMs on low-resource languages by selecting an optimal bridge language in X-ICL? Should the selection be purely data-driven, favoring high-resource bridge languages Vulic et al. (2020)? Or can human linguistic knowledge, such as language genealogy, or established evolutionary taxonomies, offer a more effective alternative Stanczak et al. (2022); Wang et al. (2024)?
To fully investigate this issue from a systematic perspective, we leverage linguistic neuron Tang et al. (2024) to guide optimal bridge language selection in X-ICL. However, there are two limitations when applying neuron-based interpretation Cao et al. (2024); Tang et al. (2024); Liu et al. (2024) on low-resource languages:
$\bullet$ Inaccurate Neuron Activation. Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal-neurons, without verifying whether LLMs truly understand the input. This may lead to unreliable neuron activations, particularly for low-resource languages. When LLMs poorly understand the probe input, they may instead activate neurons largely for processing unfamiliar or noisy input.
$\bullet$ Lacking guidance for cross-lingual transfer. Current work focuses on analyzing the distribution of neurons Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024), yet how to leverage internal neurons to enhance cross-lingual transfer remains underexplored. Recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual transfer in LLMs? This exploration is also important to transfer language neuron research to actionable strategies for enhancing LLMs’ multilinguality.
Motivated by this, we propose a simple yet effective bridge method, BridgeX-ICL, to improve LLMs’ cross-lingual capabilities, especially on low-resource languages. To address the inaccurate activation issue, we construct probe data by leveraging the ground-truth bilingual lexicon MUSE Conneau et al. (2017). We then collect bilingual word pairs that LLMs can translate accurately and prompt them to LLMs using answer generation in bi-directions. To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on HSIC Gretton et al. (2005). Furthermore, we measure the linguistic spectrum in LLMs based on overlap neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al. (2023). We conduct extensive experiments on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families. Our main contributions and findings are as follows:
  • To the best of our knowledge, this is the first work to explore language-bridge for zero-shot X-ICL to improve LLMs’ performance on low-resource languages.
  • We construct neuron probe data and use them to fully activate the anchored overlap neurons. We also propose a HSIC-based metric to quantify the similarity between overlapping-neurons and specific neurons for making optimal bridge selection in X-ICL.
  • We validate the efficacy and generalization of BridgeX-ICL on 2 cross-lingual tasks and $15$ language pairs. Here are empirical findings: 1) Strong neural overlaps align with human linguistic taxonomy within language families, but do not consistently hold across families. 2) Overlap neurons embody shared semantic information no matter among languages within or across families. 3) BridgeX-ICL improves the performance on 2 cross-lingual tasks across 15 language pairs by an average of 6.02% and 5.25% over zero-shot baselines.
Figure 1: An illustration of BridgeX-ICL approach, consisting of three steps: Neuron probe data construction; Language neurons and their overlappings detection; Optimal bridge $L_{*}$ selection based on HSIC similarity.
2 Related Work
2.1 Cross-lingual In-context Learning
LLMs face significant challenges when applied to low-resource languages Costa-jussà et al. (2022); Muennighoff et al. (2023); Huang et al. (2024), mainly due to insufficient training data and the curse of multilinguality Conneau et al. (2020). To address these issues without updating model parameters, Cross-lingual In-context Learning (X-ICL), an extension of in-context learning (ICL), has recently gained attention Brown et al. (2020a). Prior studies Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024) have demonstrated that LLMs act as effective few-shot multilingual learners, with few-shot ICL even outperforming fine-tuned language-specific models on several tasks Winata et al. (2021). However, few-shot X-ICL’s performance is highly dependent on the context and the selection of examples, especially for unconventional or ambiguous languages Philippy et al. (2023); Nazi et al. (2025). Consequently, existing research mainly focuses on optimizing few-shot example selection. To the best of our knowledge, we are the first to explore X-ICL explicitly from the perspective of leveraging language bridges.
2.2 Linguistic Neuron in LLMs
Recent research Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024); Cao et al. (2024); Wang et al. (2024) has revealed that language-related neurons exist in FFN layers of transformer architecture. Deactivating these neurons will play a vital impact on LLMs’ multilingual capacities. Beyond uncovering multilingual mechanisms, some research has gone to explore the neuron pattern across languages Wang et al. (2024); Stanczak et al. (2022) and its impact on cross-lingual performance Mondal et al. (2025); Zhang et al. (2025). Specifically, Wang et al. (2024) observed that similar languages may not exhibit significant neuron sharing in LLMs like BLOOM, suggesting that neuron sharing does not fully align with language similarity. Furthermore, recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual performance in LLMs. Motivated by these findings, we aim to further investigate LLM-internal neuron sharing across languages and its impact. In particular, we define a subset of language-overlapping neurons and explore whether they can serve as internal bridges to support cross-lingual inference.
3 Methodology
3.1 Task Statement
Given language set $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy learned by LLMs from language-overlapping neurons and then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
Figure 1 depicts three main steps of our approach: ① Neuron probe data construction; ② Language neurons and their overlappings detection; ③ Bridge language selection based on a modified HSIC dependency estimation to measure linguistic distance, selecting $L_{*}$ from the candidate set $L_{\text{candidate}}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
3.2 Probe Data Construction
We utilize two types of probe data for language neuron identification and optimal bridge selection. The former focuses on language-specific neurons and can use existing multilingual corpora. In our work, we adopt FLORES+ NLLB Team et al. (2024), a high-quality parallel corpus released by Meta, and combine its devtest and test sets to obtain 2,000 parallel sentences for each language.
Bridge selection for X-ICL needs to consider both language-specific neurons and those involved in cross-lingual tasks. Inspired by findings in task-specific neurons Song et al. (2024), we hypothesize certain neurons are associated with cross-lingual transfer whose manipulation and measurement should not rely on feeding monolingual input. Therefore, we construct probe data by leveraging bilingual word translations.
Prompt Design We collect $d$ (i.e., $100$) word pairs that LLMs can translate accurately per language pair. We prompt these word pairs to LLMs in both directions of $L_{1}$-$L_{2}$ and $L_{2}$-$L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated. Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate and pronounced neuron activation. Examples of probe data for $3$ language pairs are shown below.
3.3 Linguistic Overlap Neurons
3.3.1 Neurons in LLMs
Neurons identification is based on Tang et al. (2024), which hypothesizes that language neurons are mainly located in the Feed-Forward Network (FFN) layers. Given the transformation in the $i$-th layer:
$\boldsymbol{h}_{i}=\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})\cdot\boldsymbol{W}_{2}^{i}$ (1)
where $\tilde{\boldsymbol{h}_{i}}$ is the hidden state input to the $i$-th layer and $\sigma(\cdot)$ denotes the activation function. $\boldsymbol{W}_{1}^{i}\in\mathbb{R}^{d\times N}$ and $\boldsymbol{W}_{2}^{i}\in\mathbb{R}^{N\times d}$ are the learned parameters. Here, a neuron is defined as a linear transformation of a single column in $\boldsymbol{W}_{1}^{i}$ and there are $N$ neurons in each layer. The activation value of the $j$-th neuron is $\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})_{j}$. If this value exceeds 0, the neuron is considered activated.
3.3.2 Overlap Neuron Identification
First, we identify neurons $\mathcal{T}_{k}$ associated with specific languages $L_{k}$. Unlike existing work Wang et al. (2024); Mondal et al. (2025) using LAPE Tang et al. (2024) to identify neurons with high activation probability for one language but low for others, which is less effective to detect neuron relationships across languages, we identify neurons $\mathcal{T}_{k}$ for each language $L_{k}$ based their activation frequency. Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens of language $L_{k}$. Neurons with the top $\tau\cdot N$ activation frequencies are selected into $\mathcal{T}_{k}$ based on a threshold $\tau$.
Overlap Neuron Definition. For languages $L_{u}$, $L_{v}$, given their associated neurons $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$. For the $i$-th layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
Linguistic Neuron Similarity. We measure the linguistic similarity between $L_{u}$ and $L_{v}$ based on their overlap neurons’ activation frequency as:
$\text{sim}(\mathcal{T}_{u},\mathcal{T}_{v})=\frac{\boldsymbol{f_{u}}\cdot\boldsymbol{f_{v}}}{\lVert\boldsymbol{f_{u}}\rVert\lVert\boldsymbol{f_{v}}\rVert}$ (2)
where $f_{u,j}\in\boldsymbol{f_{u}}$ and $f_{v,j}\in\boldsymbol{f_{v}}$ denote the activation frequency of the $j$-th neuron in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$ and $L_{v}$, respectively.
3.3.3 Overlap Neurons Patterns
Second, we use the constructed probe data to explore overlap neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection. We make two observations:
Figure 2: Language-overlapping neurons on distant pair (Arabic-Swahili) and close pair (Arabic-Hebrew).
Figure 3: Layer-wise latent embeddings projected with MDS in French-Swahili translation. A rainbow-colored path traces the latent embeddings across 32 layers. The predicted Swahili tokens are in green and their correct English tokens in orange.
$\bullet$ Similar languages share more neurons than distant ones. For example, Arabic-Hebrew within the same language family has more overlapping-neurons than Arabic-Swahili across families, presented in Figure 2, This suggests the potential of neural overlap to measure language distance.
$\bullet$ Overlap neurons concentrate in middle-layers and final-layers and serve distinct roles of semantic understanding and language coding for next-token prediction. This is also evidenced by neurons deactivation in Figure 7. Neurons in final layers are task-related and are responsible for cross-lingual generation. To examine whether middle-neurons handle semantic understanding, we employ a technique called logit lens Nostalgebraist (2020) to visualize the latent embeddings passing through neuron layers. We test in French-Swahili translation and select 60 word pairs that LLaMA 3 can translate accurately. After recording the model’s latent embeddings at each layer for next-token prediction, we use classical multidimensional scaling (MDS) to embed them in a 2D space, presented in Figure 3. The embedding trajectory is marked in a rainbow-colored path (e.g., red = layers 1-4, violet = layers 25-32). We can observe French inputs and the correct English next tokens cluster in middle layers, suggesting LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning. Neurons in middle-layers should be prioritized over final-layers when measuring language similarity.
3.4 Bridge Language Selection
Based on the above observations, we quantify activation similarity between source-target overlap neurons and bridge-specific neurons and identify the optimal bridge language to facilitate X-ICL.
Given a language pair $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ when probing LLMs with $d$ samples in both directions for balanced activation in $L_{s}$ and $L_{t}$. Simultaneously, we obtain $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$, the activation value matrix for bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$, and $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents language-specific neurons in $L_{y}$ excludes neurons in $\mathcal{T}_{s,t}$ and $\mathcal{T}_{y^{\prime}}$, where $L_{y^{\prime}}\neq L_{y}$.
We employ the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al. (2005) to measure the nonlinear dependency between activation matrices of $\mathbf{X}$ and $\mathbf{Y}$. Average pooling will be performed to standardize matrices of $\mathbf{X}$ and $\mathbf{Y}$ to have the same row dimension $n$. The formal HSIC is calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$. $\mathbf{H}=\mathbf{I}_{n\times n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is a centering matrix, where $\mathbf{I}_{n\times n}$ is the identity matrix of size $n\times n$, $\mathbf{1}_{n}$ is a vector of $n$ ones. Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy to measure the strongest dependency between individual neurons in one set and the entire distribution of the other, computed as:
$\mathrm{HSIC}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=$ (3)
$\frac{1}{2}\left(\max_{i}\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\right)$
where $\mathbf{x}_{i}\in\mathbf{X},\mathbf{y}_{j}\in\mathbf{Y}$. We compute the dependency scores per layer, then average them across middle $K$ layers to get $L_{y}$’s selection probability:
$p(L_{y}|L_{s}\text{-}L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{HSIC}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
where $K$ is determined according to embedding semantic similarity and discussed in section B.3. Finally, the optimal bridge $L^{*}$ is selected by:
$L^{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\text{-}L_{t})$ (5)
4 Experiment
4.1 Experiment Setup
Implementation. We evaluate BridgeX-ICL on $2$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
The evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages. Therefore, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw), and high-to-low (e.g., En-He) pairs, both within and across families. The classification of high-, moderate-, and low-resource languages is based on their proportion in LLM’s training corpora, following previous work Cieri et al. (2016). Since the bridge language should be well supported by LLMs, we take 6 languages in Indo-European family as candidate bridges.
Datasets. To evaluate the generalization of bridge selection beyond the BLI task, we also evaluate on an additional cross-lingual task: Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al. (2024).
To evaluate low-resource languages, a key challenge lies in lacking evaluation benchmarks. Although the ground truth MUSE Conneau et al. (2017) provides $110$ bilingual dictionaries for BLI task, it does not cover the tested 15 language pairs. To solve this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English. For languages not in MUSE (e.g., Swahili), we extracted word pairs from wiktionary_bli Izbicki (2022) to build En-Sw. We checked all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems. The constructed BLI dictionaries are available at: https://anonymous.4open.science/r/BLI--0481/.
Metrics. For BLI task, we use Precision@N metric, which measures the accuracy of the model’s top $N$ candidate translations. In this study, $N$ is set to 1. For the MRC task, we use accuracy to measure whether the model selects the correct answer from multiple choices.
LLMs. We conducted experiments on two open-source LLMs: LLaMA-3-8B Grattafiori et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). Their training corpora cover 176 and 53 languages, respectively, which include all the experimental low-resource languages and allow us to explore the underlying linguistic mechanisms.
[…]
4.2 Main Results
4.2.1 LLMs’ Linguistic Spectrum Discussion
(a) LLaMA 3’s Linguistic spectrum
(b) Human language similarity from Glottolog
Figure 4: Comparison of linguistic spectrum calculated based on overlap neurons in LLaMA 3 and human language similarity derived from Glottolog Phylogenetic Trees, including 15 languages from 7 families. Darker blue indicates a higher language similarity.
This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlap neurons in LLaMA 3 and Mistral, as presented in Figure 4 (a) and Figure 8 in B.2, respectively. To evaluate how closely the linguistic spectrum learned by LLMs align with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al. (2023), which encode hierarchical relationships among 8,000+ human languages, to draw human language similarity in Figure 4(b). The detailed linguistic similarity computed based on Glottolog is in Appendix A. In Figure 4, color intensity represents degree of similarity between languages, with darker blue indicating stronger degree and the diagonal is self-similarity (1.0).
Table 1: Comparison of BLI task improvement on 15 language pairs. The highest gains are marked with bold in few-shot and zero-shot with bridge methods. ’-’ indicates the selected bridge is either the source or target language.
LLaMA-3-8B
Method Zh-Ja Zh-He Zh-Tl Zh-Sw Ar-Ja Ar-He Ar-Tl Ar-Sw Id-Ja Id-He Id-Tl Id-Sw En-He En-Tl En-Sw
Zero-shot 67.10 44.10 42.60 31.20 69.90 47.00 46.70 39.10 62.50 44.70 49.30 25.90 56.90 60.00 28.80
One-shot +3.20 +12.60 +1.20 +3.00 +4.20 +13.50 -4.80 +0.60 +7.40 +7.50 +0.70 +6.00 +18.60 -6.30 +4.40
Two-shot +9.40 +16.90 +2.20 +5.10 +9.40 +13.90 -0.30 +3.50 +16.00 +15.90 +5.90 +6.10 +23.90 -5.50 +6.50
Three-shot +6.70 +22.20 +3.70 +6.80 +7.80 +16.90 +1.50 +3.70 +16.70 +22.90 +10.10 -4.30 +26.30 -4.00 +6.90
Few-shot Four-shot +12.50 +20.50 +3.20 +7.00 +7.70 +15.80 +0.80 +3.20 +14.30 +22.00 +10.60 -6.00 +26.10 -3.70 +7.40
Human source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
Human target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
English bridge +10.80 +12.60 +11.40 +4.80 +10.70 +17.50 +10.10 +3.30 +3.60 +9.30 +10.40 +2.50 - - -
Sharing matter +9.50 +14.50 +8.40 +2.60 +6.40 +17.10 +6.10 +3.90 +2.60 +15.20 +7.70 +2.40 +12.30 -3.30 -1.70
IoU Score +10.80 +12.60 +11.40 +4.80 +10.70 +13.20 +5.30 +3.50 +2.30 +15.20 +7.40 +3.80 +9.70 -6.10 -1.80
LAPE_overlap +10.50 +13.90 +11.40 +3.00 +10.70 +17.50 +10.10 +3.30 +3.60 +9.30 +10.40 +2.50 +12.30 -6.10 -1.80
Zero-shot
with bridge
Ours +10.80 +12.60 +11.40 +4.80 +10.70 +17.50 +10.10 +3.30 +3.60 +16.60 +11.70 +4.10 +14.90 -1.30 -2.60
Mistral-7B
Method Zh-Ja Zh-He Zh-Tl Zh-Sw Ar-Ja Ar-He Ar-Tl Ar-Sw Id-Ja Id-He Id-Tl Id-Sw En-He En-Tl En-Sw
Zero-shot 57.80 26.20 34.10 8.40 52.50 32.30 28.00 9.10 48.40 36.20 40.60 8.60 47.80 45.70 8.20
One-shot -7.20 -0.40 -1.90 +1.60 -8.10 -7.40 -2.00 +2.10 +5.60 +2.50 -3.70 +1.20 +0.40 -13.80 +2.40
Two-shot +0.40 +0.80 -2.90 +2.10 -5.10 -7.00 -0.40 +1.60 +10.10 +3.20 -0.60 +1.60 +0.50 -3.90 +3.20
Three-shot +3.00 +0.70 -1.60 +2.20 -2.90 -6.50 +0.20 +2.00 +10.70 +3.30 +0.80 +2.00 0.00 -2.80 +3.00
Few-shot Four-shot +3.20 +1.20 -0.50 +2.60 -2.90 -6.60 0.00 +2.20 +10.70 +4.10 +2.50 +2.60 +1.20 -2.00 +2.50
Human source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
Human target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
English bridge +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 - - -
Sharing matter +7.90 +8.60 +6.40 +1.20 +4.60 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 -0.10 +1.50 +0.90
IoU Score +0.50 +5.30 +4.90 +2.20 +5.90 +1.10 +7.50 +1.60 +2.40 -2.80 -0.50 +1.00 -1.00 +1.50 +2.20
LAPE_overlap +0.50 +3.50 +2.20 +2.20 +8.90 +2.70 +7.60 +1.20 +9.60 +2.20 +1.00 +0.20 -0.10 +4.50 +2.00
Zero-shot
with bridge
Ours +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +2.40 -3.30 +3.30 +1.00 -0.10 +1.50 +2.00
Linguistic spectrum learned by LLMs are not fully aligned with human languages. We can observe a strong neural similarity within language families, marked using red text box in Figure 4, which matches human linguistic taxonomy. For example, within Afro-Asiatic family, both from Arabic (Ar) and Hebrew (He), have a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161. In addition, high-resource Indo-European languages, such as Fr-It and Pt-Es, show the highest overlap scores, with darkest blue in the bottom-right corner of the heatmap. But this alignment breaks down between high- and low-resource languages, like Arabic-French with similarity $0.479$.
LLMs build their own distinct understanding of language relationships. The calculated linguistic spectrum of LLaMA 3 and Mistral are similar but not the same. The two models may choose different bridges for the same language pair, as discussed later. This is likely because their internal linguistic relationships are primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al. (2023). This explains why Arabic has the strongest similarity ($0.479$) with French in Romance family, rather than with Hebrew ($0.470$) in the same Afro-Asiatic.
4.2.2 Cross-lingual Results Analysis
This section compares performance of BridgeX-ICL against various baselines on BLI task across $15$ language pairs, as presented in Table 1.
Table 2: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selection and bold marks the highest gains at each language pair.
LLaMA-3-8B Mistral-7B
Bridge Zero-shot Ours Bridge Zero-shot Ours
Zh-Ja En 61.80 -0.40 En 66.20 +6.20
Zh-He En 56.00 +7.20 En 50.20 +10.20
Zh-Tl En 57.20 +3.00 En 60.60 +7.80
Zh-Sw En 48.60 +10.40 Pt 43.00 +1.00
Ar-Ja En 52.20 +3.20 En 48.40 +7.40
Ar-He En 51.20 +4.00 En 42.00 +5.20
Ar-Tl En 46.40 +7.20 En 47.60 +4.60
Ar-Sw En 40.60 +12.60 En 30.40 +9.80
Id-Ja En 56.20 +8.20 Pt 63.40 +2.00
Id-He Es 58.20 +7.00 Es 45.80 +6.40
Id-Tl Fr 58.40 +4.20 De 55.60 +3.00
Id-Sw Fr 49.80 +5.20 Pt 39.60 -2.80
En-He Es 70.60 +8.00 Es 61.20 +3.80
En-Tl Fr 72.80 +3.20 Es 72.20 +2.40
En-Sw Fr 64.60 +7.40 Fr 51.20 +0.20
We can find: 1) LLMs exhibit poor and imbalanced performance on low-resource languages. For example, LLaMA 3 achieves its best BLI performance of $69.90$ on Ar-Ja pair, but worst of $25.90$ on Id-Sw pair. 2) LLMs are few-shot multilingual learners. However, few-shot X-ICL does not consistently yield stable gains, e.g., one-shot falling below zero-shot, and will not lead to further gains when the number of shots goes beyond $3$. It indicates when applying few-shot X-ICL, 3-shots will be enough. 3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages. BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method. While human source/target methods are the least effective. It seems using English as the default bridge is cost-effective, which will be discussed in section 5.1.
5 Discussion
5.1 Application of Bridge Language
Beyond BLI task, Table 2 evaluates BridgeX-ICL on the MRC cross-lingual task. The prompt for zero-shot with bridge in MRC is detailed in Appendix E. Results show our approach still works well on MRC task and benefits more on LLaMA 3 than Mistral. For example, BridgeX-ICL improves performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across 15 language pairs, while the average improvement is 4.484.48% on Mistral.
English is selected as the optimal bridge in 9 out of 15 language pairs. The reason is partly due to LLMs’ unbalanced language abilities across 5 candidate bridges. As discussed in Figure 3 another key factor is LLMs’s inherent preference of English-pivot during cross-lingual transfer.
5.2 Ablation Study
In this part, we conduct ablation study to evaluate the impact of neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task for example.
Table 3: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from FLORES+ dataset.
LLaMA-3-8B Mistral-7B
w/o * Ours w/o * Ours
Zh-Ja 76.40 77.90 $\uparrow$ 63.40 65.70 $\uparrow$
Zh-He 56.70 56.70  32.70 34.80 $\uparrow$
Zh-Tl 51.10 54.00 $\uparrow$ 38.20 40.50 $\uparrow$
Zh-Sw 36.40 36.00 $\downarrow$ 10.60 9.60 $\downarrow$
Ar-Ja 77.80 80.60 $\uparrow$ 58.40 61.40 $\uparrow$
Ar-He 64.10 64.50 $\uparrow$ 33.40 35.00 $\uparrow$
Ar-Tl 52.00 56.80 $\uparrow$ 33.30 35.50 $\uparrow$
Ar-Sw 41.60 42.40 $\uparrow$ 10.70 10.70
Id-Ja 65.10 66.10 $\uparrow$ 50.80 58.00 $\uparrow$
Id-He 59.90 61.30 $\uparrow$ 33.40 38.40 $\uparrow$
Id-Tl 57.00 61.00 $\uparrow$ 40.20 40.20
Id-Sw 28.30 30.00 $\uparrow$ 9.20 9.20
En-He 66.60 71.80 $\uparrow$ 47.70 47.70
En-Tl 58.70 58.70  47.70 48.00 $\uparrow$
En-Sw 26.20 26.20  10.20 10.30 $\uparrow$
Table 3 presents the results of ablation experiments to compare “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al. (2024). Appendix C illustrates the detailed construction for “w/o *”. These results highlight impact of probing data in neuron manipulation. Table 4 in Appdedix C compares the performance of HSIC and Cosine similarity and shows the proposed HSIC helps to capture the dependency between language-overlapping neurons and specific neurons.
5.3 Overlapping Neuron Distribution
Figure 5: Distribution of overlap neurons in language pairs within and across families in LLaMA 3.
Figure 6: Layer-wise activation frequency of neurons viewed from language families in LLaMA 3.
This section analyzes the distribution of overlap neurons. Figure 5 compares neurons in language pairs within the same family (e.g., Ar-He) and across families (e.g., Ar-Sw). Obviously, Ar-He shares more overlap neurons. Similar observations can be found in comparing language pairs from different source languages (Zh, Ar, Id, and En ) to a same target language He (Figure 10 in B.4) From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages. Obviously, low-resource languages within the Uralic family have the highest activation frequency, while Indo-European languages have the lowest. We hypothesize LLMs activate neurons more frequently for processing low-resource languages due to their perceived difficulty.
6 Conclusion
In this work, we explore whether sharing neurons can improve LLMs’ cross-lingual performance on low-resource languages. We propose a simple yet effective language-bridge approach with the help of neuron interpretation. To ensure accurately and fully activate overlap neurons across languages, we construct neurons probe data from the ground-truth MUSE dictionaries. By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate the efficacy and generalization of our approach.
Limitations
This work focuses on sharing neurons across languages and relies on evaluated data to validate the effectiveness of our approach. Due to the limitation of lacking evaluation benchmarks for low-resource languages, the experiments are conducted on 15 language pairs and have not been extensively validated on a wide range of low-resource languages. We select 4 typologically diverse low-resource languages from distinct families and test their performance on 2 cross-lingual tasks. Second, our study reveals that high-quality probe data is essential for accurately analyzing neurons’ behaviors of low-resource languages. The proposed linguistic distance measurement is probe-data-induced, providing qualitative insights, but lacks quantitative precision.
Acknowledgments
This work was supported by the National Social Science Foundation (No.24CYY107) and the Fundamental Research Funds for the Central Universities (No.2024TD001).
Appendix A Appendix: Linguistic Similarity Based on Glottolog Phylogenetic Trees
We leverage Glottolog verion 5.1 Hammarström et al. (2023) as a foundational phylogenetic framework to calculate the linguistic similarity of human languages. It has two key steps: data preprocessing and similarity calculation.
Data Preprocessing. The preprocessing pipeline consists of three steps: 1) Glottocode Identifiers with regex pattern matching to ensure unambiguous language node identification. 2) Standardize node naming with underscores (e.g., [sini1245]→ _sini1245_), ensuring consistent formatting in downstream phylogenetic analyses. 3) Mitigate encoding conflicts through temporary file caching. These steps preserve both accurate phylogenetic tree parsing and computational compatibility.
Similarity Calculation. The proposed metric integrates two well-established phylogenetic principles from historical linguistics, which includes node distance normalization and depth-adjusted compensation.
First, building upon Wichmann & Holman’s framework for typological stability assessment Wichmann and Holman (2009), we compute the inter-language distance $d(L_{1},L_{2})$ between languages $L_{1}$ and $L_{2}$ using ETE3’s optimized tree traversal algorithms Huerta-Cepas et al. (2016) and then normalize it to make distances comparable across language families, calculated as:
[…]
where $\hat{D}$ is the family-specific maximum. For example, $\hat{D}=80$ for Sino-Tibetan languages, reflecting their deep internal divergence, whereas $\hat{D}=75$ for Indo-European languages, due to their relatively shallower subgroup structure.
Second, depth-adjusted compensation aims to mitigate biases introduced by uneven tree depth and family-specific structural variation. Following the work Gray et al. (2009) to calculate depth disparity factor $\delta(L_{1},L_{2})$, we measure the depth $\alpha_{\text{depth}}(L_{1},L_{2})$ between $L_{1}$ and $L_{2}$ as:
$\alpha_{\mathrm{depth}}=1-\frac{\delta(L_{1},L_{2})}{\max(\text{depth}(L_{1}),\text{depth}(L_{2}))}$ (7)
The final language similarity score is computed as:
$\mathrm{Sim}(L_{1},L_{2})=S_{\mathrm{distance}}\times\alpha_{\mathrm{depth}}$ (8)
Appendix B Appendix: Neuron Patterns
B.1 Deactivation Overlap Neurons
Figure 7 presents overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
Figure 7: Overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
B.2 Linguistic Spectrum in Mistral
Figure 8: Mistral’s linguistic spectrum across 15 languages from 7 families. The color intensity represents the degree of overlap between language pairs.
B.3 Parameter $K$ Discussion
Here we discuss which $k$ middle layers should be selected to quantify linguistic similarity. According to observations in section 3.3.3, neurons in middle-layers should be prioritized over final-layers when measuring language similarity. We use embedding semantic similarity metric to determine $K$. We compute the layer-wise embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew pairs when predicting the same Hebrew token. As presented in Figure 9, we find embedding similarity is stable in the middle layers 10-21 and is not sensitive to variations in the predicted tokens. Therefore, $K$ layers is set to be 10-21 in LLaMA 3 and 15-23 in Mistral.
Figure 9: Embedding semantic similarity between Arabic-to-Hebrew and Chinese-to-Hebrew translations when predicting the same token at each layer.
B.4 Overlap Neuron Distribution
Figure 10 presents the distribution of overlap neurons across different language pairs in LLaMA 3.
Figure 10: Distribution of overlap neurons across different language pairs in LLaMA 3.
Appendix C Appendix: Ablation Results
This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on bridge selection. Table 3 and Table 4 present the results of ablation experiments on the BLI task. “w/o *” denotes replacing our probe data with a simplified version based on FLORES+. For example, “w/o *” probe data in Indonesian-Hebrew is illustrated in Figure 11.
Table 4: Performance comparison of using HSIC and Cosine similarity metrics on the BLI task.
LLaMA-3-8B
Bridge HSIC Bridge Cosine
Zh-Ja En 77.90 $\uparrow$ De 76.60
Zh-He En 56.70 $\downarrow$ De 58.60
Zh-Tl En 54.00 $\uparrow$ De 51.00
Zh-Sw En 36.00 $\uparrow$ De 33.80
Ar-Ja En 80.60 $\uparrow$ De 76.30
Ar-He En 64.50 $\uparrow$ De 64.10
Ar-Tl En 56.80 $\uparrow$ De 52.80
Ar-Sw En 42.40 $\downarrow$ De 43.00
Id-Ja En 66.10 $\uparrow$ De 65.10
Id-He Es 61.30 $\uparrow$ De 59.90
Id-Tl Fr 61.00 $\uparrow$ De 57.00
Id-Sw Fr 30.00 $\uparrow$ De 28.30
En-He Es 71.80 $\downarrow$ Pt 72.30
En-Tl Fr 58.70 $\uparrow$ Pt 55.80
En-Sw Fr 26.20 $\uparrow$ It 21.50
Figure 11: Example of “w/o *” probe data in Indonesian-Tagalog.
Appendix D Appendix: Neuron Semantic Analysis
In this section, we analyze the semantic similarity of overlapping neurons in two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, different language family) and Portuguese-Spanish-Italian (pt-Es-It, same language families). For each group, we select overlapping neurons and the same number of randomly sampled neurons for comparison. We input $m$ parallel sentences and record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for overlap and random neurons. These matrices are mapped to a 2D semantic space using UMAP McInnes et al. (2018), with each point representing a neuron activated by a sentence.
As shown in Figure 12, randomly sampled neurons align with linguistic relationships. For example, random neurons of He and Tl are close, indicating semantic proximity in the same family. The overlap neurons cluster together, both within and across language families, proving that our approach can effectively capture semantic similarity.
Figure 12: Visualization of semantic similarity comparing language-overlapping neurons and randomly sampled neurons.
Appendix E Appendix: Prompt Templates

### v2

@@
 Abstract
-The current Large Language Models (LLMs) face significant challenges in improving performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning.
-From the perspective of language-bridge, we propose BridgeX-ICL, a simple yet effective method to improve zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages.
-Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs or not.
-We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly, to ensure full activation of these anchored neurons.
-Subsequently, we propose an HSIC-based metric to quantify LLMs-internal linguistic spectrum based on overlap neurons, which guides optimal bridge selection.
-The experiments conducted on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families (covering both high-low and moderate-low pairs) validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs.
+The current Large Language Models (LLMs) face significant challenges in improving their performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning.
+From the perspective of language-bridge, we propose a simple yet effective method, namely BridgeX-ICL, to improve the zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages.
+Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs.
+We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly to ensure full activation of these anchored neurons.
+Subsequently, we propose an HSIC-based metric to quantify LLMs’ internal linguistic spectrum based on overlapping neurons, guiding optimal bridge selection.
+The experiments conducted on $4$ cross-lingual tasks and $15$ language pairs from $7$ diverse families, covering both high-low and moderate-low pairs, validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs.
+The code is publicly available at https://github.com/xuyuemei/BridgeX-ICL.
+Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages
+Yuemei Xu1, Kexin Xu1, Jian Zhou1, Ling Hu1, Lin Gui2 1 School of Information Science and Technology, Beijing Foreign Studies University 2 Department of Informatics, King’s College London {xuyuemei, xukexin, bwzj, huling}@bfsu.edu.cn lin.1.gui@kcl.ac.uk
 1 Introduction
+footnotetext: This work was supported by the National Social Science Foundation (No.24CYY107), the Fundamental Research Funds for the Central Universities (No.2024TD001), and the National Natural Science Foundation of China (No.
+62576120).
 Although Large Language Models (LLMs) have demonstrated impressive multilingual capacities, there is still significant space for improving the performance on low-resource languages Huang et al.
-(2024); Nazi et al.
+(2024); Al Nazi et al.
 (2024).
@@
 (2021); Tanwar et al.
-(2023); Nazi et al.
+(2023); Al Nazi et al.
 (2024); Cahyawijaya et al.
-(2024), which surprisingly works well on low-resource languages, primarily due to LLMs are in-context low-resource language learners Brown et al.
+(2024), which surprisingly works well on low-resource languages, likely because LLMs are in-context low-resource language learners Brown et al.
 (2020b); Cahyawijaya et al.
@@
 For instance, in the Arabic-to-Hebrew Bilingual Lexicon Induction (BLI) task, the zero-shot baseline accuracy in LLaMA 3 is 47.0%.
-However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which significantly outperforms even the two-shot X-ICL.
+However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which even significantly outperforms the two-shot X-ICL.
 This observation motivates us to further explore: How can we improve cross-lingual capabilities of LLMs on low-resource languages by selecting an optimal bridge language in X-ICL?
@@
 (2024)?
-To fully investigate this issue from a systematic perspective, we leverage linguistic neuron Tang et al.
-(2024) to guide optimal bridge language selection in X-ICL.
+To systematically investigate this issue, we leverage linguistic neurons Tang et al.
+(2024) that handle language features to guide optimal bridge language selection in X-ICL.
 However, there are two limitations when applying neuron-based interpretation Cao et al.
@@
 $\bullet$ Inaccurate neuron activation.
-Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal-neurons, without verifying whether LLMs truly understand the input.
+Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal neurons, without verifying whether LLMs truly understand the multilingual input.
 This may lead to unreliable neuron activations, particularly for low-resource languages.
-When LLMs poorly understand the probe input, they may instead activate neurons largely for processing unfamiliar or noisy input.
+When LLMs poorly understand the probe input, they may instead activate neurons for processing unfamiliar or noisy input.
 $\bullet$ Lacking guidance for cross-lingual transfer.
-Current work focuses on analyzing the distribution of neurons Stanczak et al.
-(2022); Tang et al.
-(2024); Liu et al.
-(2024), yet how to leverage internal neurons to enhance cross-lingual transfer remains underexplored.
 Recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al.
 (2025).
-This raises a critical question: Whether sharing neurons can improve cross-lingual transfer in LLMs?
-This exploration is also important to transfer language neuron research to actionable strategies for enhancing LLMs’ multilinguality.
+This raises a critical question: if language-specific neurons cannot, can sharing neurons improve cross-lingual transfer in LLMs?
+This exploration is also important for transferring language neuron research to actionable strategies to enhance the multilinguality of LLMs.
 Motivated by this, we propose a simple yet effective bridge method, BridgeX-ICL, to improve LLMs’ cross-lingual capabilities, especially on low-resource languages.
@@
 (2017).
-We then collect bilingual word pairs that LLMs can translate accurately and prompt them to LLMs using answer generation in bi-directions.
-To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on HSIC Gretton et al.
+We collect bilingual word pairs from MUSE that LLMs can translate accurately and use them to prompt the models for bidirectional translations, generating answers in both language directions.
+To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al.
 (2005).
-Furthermore, we measure the linguistic spectrum in LLMs based on overlap neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al.
+Furthermore, we measure the linguistic spectrum in LLMs based on overlapping-neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al.
 (2023).
-We conduct extensive experiments on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families.
+We conduct extensive experiments on $4$ cross-lingual tasks and $15$ language pairs from $7$ diverse families.
 Our main contributions and findings are as follows:
   • To the best of our knowledge, this is the first work to explore language-bridge for zero-shot X-ICL to improve LLMs’ performance on low-resource languages.
-  • We construct neuron probe data and use them to fully activate the anchored overlap neurons.
-We also propose a HSIC-based metric to quantify the similarity between overlapping-neurons and specific neurons for making optimal bridge selection in X-ICL.
-  • We validate the efficacy and generalization of BridgeX-ICL on 2 cross-lingual tasks and $15$ language pairs.
+  • We construct accurate neuron probe data and use it to fully activate the anchored overlap neurons.
+We also propose an HSIC-based metric to quantify the similarity between overlapping neurons and specific neurons for optimal bridge selection in X-ICL.
+  • We validate the generalization of BridgeX-ICL on $4$ cross-lingual tasks and $15$ language pairs.
 Here are empirical findings: 1) Strong neural overlaps align with human linguistic taxonomy within language families, but do not consistently hold across families.
-2) Overlap neurons embody shared semantic information no matter among languages within or across families.
-3) BridgeX-ICL improves the performance on 2 cross-lingual tasks across 15 language pairs by an average of 6.02% and 5.25% over zero-shot baselines.
+2) Overlapping neurons embody shared semantic information, regardless of the language within or between families.
+3) BridgeX-ICL improves the performance on cross-lingual tasks of BLI and MRC across 15 language pairs by an average of $6.02\%$ and $5.25\%$ over zero-shot baselines.
+4) English is selected as the optimal bridge in 9 out of 15 language pairs in LLaMA (7 out of 15 in Mistral), indicating that high-resource, Latin-script languages tend to be the default bridge.
+We also find that non-Latin script languages like Chinese also show potential as effective bridges.
 Figure 1: An illustration of BridgeX-ICL approach, consisting of three steps: Neuron probe data construction; Language neurons and their overlappings detection; Optimal bridge $L_{*}$ selection based on HSIC similarity.
@@
 (2021); Tanwar et al.
-(2023); Nazi et al.
+(2023); Al Nazi et al.
 (2024); Cahyawijaya et al.
@@
 However, few-shot X-ICL’s performance is highly dependent on the context and the selection of examples, especially for unconventional or ambiguous languages Philippy et al.
-(2023); Nazi et al.
+(2023); Al Nazi et al.
 (2024).
@@
 (2024) has revealed that language-related neurons exist in FFN layers of transformer architecture.
-Deactivating these neurons will play a vital impact on LLMs’ multilingual capacities.
+Deactivating these neurons will have a vital impact on LLMs’ multilingual capacities.
 Beyond uncovering multilingual mechanisms, some research has gone to explore the neuron pattern across languages Wang et al.
@@
 3.1 Task Statement
-Given language set $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy learned by LLMs from language-overlapping neurons and then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
-Figure 1 depicts three main steps of our approach: ① Neuron probe data construction; ② Language neurons and their overlappings detection; ③ Bridge language selection based on a modified HSIC dependency estimation to measure linguistic distance, selecting $L_{*}$ from the candidate set $L_{\text{candidate}}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
+Given a set of languages $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy implicitly learned by LLMs from language overlapping neurons, then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
+Figure 1 depicts three main steps of our approach: ① neuron probe data construction; ② language neurons and their overlapping detection; ③ bridge language selection, guided by the observed pattern of overlapping neurons and a modified HSIC dependency estimation, which selects the optimal $L_{*}$ from the candidate set $\mathcal{L}_{\text{candidate}}$ to facilitate X-ICL from a source language $L_{s}$ to a target language $L_{t}$.
 3.2 Probe Data Construction
-We utilize two types of probe data for language neuron identification and optimal bridge selection.
-The former focuses on language-specific neurons and can use existing multilingual corpora.
+We employ two types of probe data for different purposes of language neurons identification and optimal bridge selection.
+The former is task independent, targeting language neurons, and can use existing multilingual corpora.
 In our work, we adopt FLORES+ NLLB Team et al.
-(2024), a high-quality parallel corpus released by Meta, and combine its devtest and test sets to obtain 2,000 parallel sentences for each language.
-Bridge selection for X-ICL needs to consider both language-specific neurons and those involved in cross-lingual tasks.
-Inspired by findings in task-specific neurons Song et al.
-(2024), we hypothesize certain neurons are associated with cross-lingual transfer whose manipulation and measurement should not rely on feeding monolingual input.
-Therefore, we construct probe data by leveraging bilingual word translations.
-Prompt Design We collect $d$ (i.e., $100$) word pairs that LLMs can translate accurately per language pair.
-We prompt these word pairs to LLMs in both directions of $L_{1}$-$L_{2}$ and $L_{2}$-$L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated.
-Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate and pronounced neuron activation.
+(2024), a high-quality parallel corpus released by Meta, and combine its development set and test set to obtain $2,000$ parallel sentences for each language.
+Bridge selection for X-ICL needs to consider both language neurons and those that contribute to cross-lingual tasks.
+Inspired by findings on task-specific neurons Song et al.
+(2024), we propose that certain neurons directly influence cross-lingual transfer, and their manipulation and measurement should not rely solely on monolingual corpora.
+Therefore, we construct probe data by leveraging bilingual word translations as follows.
+Probe Data Design.
+We collect $d$ (i.e., $100$) word pairs that LLMs can accurately translate.
+These word pairs are fed into the LLMs in both directions of $L_{1}\to L_{2}$ and $L_{2}\to L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated.
+Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate neuron activation.
 Examples of probe data for $3$ language pairs are shown below.
@@
 3.3.1 Neurons in LLMs
-Neurons identification is based on Tang et al.
-(2024), which hypothesizes that language neurons are mainly located in the Feed-Forward Network (FFN) layers.
-Given the transformation in the $i$-th layer:
+Neuron identification follows Tang et al. (2024), which assumes that language neurons are mainly located in the Feed-Forward Network (FFN) layers. Given the transformation at the $i$-th layer:
 $\boldsymbol{h}_{i}=\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})\cdot\boldsymbol{W}_{2}^{i}$ (1)
@@
 The activation value of the $j$-th neuron is $\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})_{j}$.
-If this value exceeds 0, the neuron is considered activated.
+If this value exceeds 0, the neuron is considered an activated neuron.
 3.3.2 Overlap Neuron Identification
-First, we identify neurons $\mathcal{T}_{k}$ associated with specific languages $L_{k}$.
+First, we identify neurons $\mathcal{T}_{k}$ associated with each language $L_{k}$.
 Unlike existing work Wang et al.
 (2024); Mondal et al.
-(2025) using LAPE Tang et al.
-(2024) to identify neurons with high activation probability for one language but low for others, which is less effective to detect neuron relationships across languages, we identify neurons $\mathcal{T}_{k}$ for each language $L_{k}$ based their activation frequency.
-Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens of language $L_{k}$.
+(2025) using Language Activation Probability Entropy (LAPE) Tang et al.
+(2024) to identify neurons with high activation probability for one language but low for others, which is less effective to capture neuron relationships across languages, we identify neurons set $\mathcal{T}_{k}$ based on activation frequency.
+Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens from language $L_{k}$.
 Neurons with the top $\tau\cdot N$ activation frequencies are selected into $\mathcal{T}_{k}$ based on a threshold $\tau$.
 Overlap Neuron Definition.
-For languages $L_{u}$, $L_{v}$, given their associated neurons $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$.
-For the $i$-th layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
-Linguistic Neuron Similarity. We measure the linguistic similarity between $L_{u}$ and $L_{v}$ based on their overlap neurons’ activation frequency as:
+For languages $L_{u}$ and $L_{v}$, with associated neuron sets $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are defined as the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$.
+At the $i$-th FFN layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
+Linguistic Similarity Calculation.
+The linguistic similarity between $L_{u}$ and $L_{v}$ is quantified through the activation frequencies of their overlapping neurons.
+Let $\boldsymbol{f_{u}}=\{f_{u,1},f_{u,2},...,f_{u,|\mathcal{T}_{u,v}|}\}$ denote the activation frequency vector of neurons in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$.
+The linguistic similarity between $L_{u}$ and $L_{v}$ is calculated:
 $\text{sim}(\mathcal{T}_{u},\mathcal{T}_{v})=\frac{\boldsymbol{f_{u}}\cdot\boldsymbol{f_{v}}}{\lVert\boldsymbol{f_{u}}\rVert\lVert\boldsymbol{f_{v}}\rVert}$ (2)
-where $f_{u,j}\in\boldsymbol{f_{u}}$ and $f_{v,j}\in\boldsymbol{f_{v}}$ denote the activation frequency of the $j$-th neuron in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$ and $L_{v}$, respectively.
-3.3.3 Overlap Neurons Patterns
-Second, we use the constructed probe data to explore overlap neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection.
-We make two observations:
+By computing pairwise similarities for all languages in $\mathcal{L}$, we obtain a comprehensive linguistic spectrum of the model.
+3.3.3 Overlap Neuron Pattern
+Second, we use the constructed probe data from FLORES+ to explore overlapping neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection.
+We have two observations:
 Figure 2: Language-overlapping neurons on distant pair (Arabic-Swahili) and close pair (Arabic-Hebrew).
@@
 A rainbow-colored path traces the latent embeddings across 32 layers.
-The predicted Swahili tokens are in green and their correct English tokens in orange.
+The predicted Swahili tokens are in green, and their correct English tokens are in orange.
 $\bullet$ Similar languages share more neurons than distant ones.
-For example, Arabic-Hebrew within the same language family has more overlapping-neurons than Arabic-Swahili across families, presented in Figure 2, This suggests the potential of neural overlap to measure language distance.
-$\bullet$ Overlap neurons concentrate in middle-layers and final-layers and serve distinct roles of semantic understanding and language coding for next-token prediction.
-This is also evidenced by neurons deactivation in Figure 7.
-Neurons in final layers are task-related and are responsible for cross-lingual generation.
-To examine whether middle-neurons handle semantic understanding, we employ a technique called logit lens Nostalgebraist (2020) to visualize the latent embeddings passing through neuron layers.
-We test in French-Swahili translation and select 60 word pairs that LLaMA 3 can translate accurately.
-After recording the model’s latent embeddings at each layer for next-token prediction, we use classical multidimensional scaling (MDS) to embed them in a 2D space, presented in Figure 3.
+For example, Arabic-Hebrew within the same language family has more overlapping neurons than Arabic-Swahili across families, as presented in Figure 2.
+This suggests the potential of neural overlap to measure language distance.
+$\bullet$ Overlap neurons are predominantly concentrated in the middle and final layers, serving distinct roles of semantic understanding and language decoding for next-token prediction.
+This neural function is further evidenced by neurons deactivation shown in Figure 7.
+Specifically, neurons in final layers are task-related and responsible for cross-lingual generation.
+To examine whether middle-layer neurons handle semantic understanding, we further employ a technique called logit lens Nostalgebraist (2020) to visualize the latent semantic embeddings across layers.
+We visualize French-Swahili translation using $60$ word pairs that LLaMA 3 translates accurately.
+We extract the model’s latent embeddings at each layer for next-token prediction and project them into a 2D space using classical multidimensional scaling (MDS), presented in Figure 3.
 The embedding trajectory is marked in a rainbow-colored path (e.g., red = layers 1-4, violet = layers 25-32).
-We can observe French inputs and the correct English next tokens cluster in middle layers, suggesting LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning.
-Neurons in middle-layers should be prioritized over final-layers when measuring language similarity.
+We can observe French inputs and their corresponding correct English next tokens cluster in middle layers, indicating that LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning there.
+Neurons in middle layers should be prioritized over those in final layers when quantifying language similarity.
 3.4 Bridge Language Selection
-Based on the above observations, we quantify activation similarity between source-target overlap neurons and bridge-specific neurons and identify the optimal bridge language to facilitate X-ICL.
-Given a language pair $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ when probing LLMs with $d$ samples in both directions for balanced activation in $L_{s}$ and $L_{t}$.
-Simultaneously, we obtain $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$, the activation value matrix for bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$, and $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents language-specific neurons in $L_{y}$ excludes neurons in $\mathcal{T}_{s,t}$ and $\mathcal{T}_{y^{\prime}}$, where $L_{y^{\prime}}\neq L_{y}$.
-We employ the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al.
-(2005) to measure the nonlinear dependency between activation matrices of $\mathbf{X}$ and $\mathbf{Y}$.
+Based on the above observations, we leverage the constructed probe data with $d$ samples per language pair to identify the optimal bridge language $L_{*}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
+Given $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$ identified in section 3.3.2, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ by prompting the LLM with $d$ samples in both directions for balanced neuron activation in $L_{s}$ and $L_{t}$.
+We also obtain the activation matrix $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$ for a candidate bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$.
+Here, $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents the set of language-specific neurons in $L_{y}$, excluding those shared with $\mathcal{T}_{s,t}$ or with any other language $L_{y^{\prime}}\neq L_{y}$.
+We employ HSIC Gretton et al.
+(2005) to measure the nonlinear dependency between activation matrices $\mathbf{X}$ and $\mathbf{Y}$.
 Average pooling will be performed to standardize matrices of $\mathbf{X}$ and $\mathbf{Y}$ to have the same row dimension $n$.
-The formal HSIC is calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$.
+The HSIC is formally calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$.
 $\mathbf{H}=\mathbf{I}_{n\times n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is a centering matrix, where $\mathbf{I}_{n\times n}$ is the identity matrix of size $n\times n$, $\mathbf{1}_{n}$ is a vector of $n$ ones.
-Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy to measure the strongest dependency between individual neurons in one set and the entire distribution of the other, computed as:
-$\mathrm{HSIC}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=$ (3)
-$\frac{1}{2}\left(\max_{i}\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\right)$
-where $\mathbf{x}_{i}\in\mathbf{X},\mathbf{y}_{j}\in\mathbf{Y}$. We compute the dependency scores per layer, then average them across middle $K$ layers to get $L_{y}$’s selection probability:
-$p(L_{y}|L_{s}\text{-}L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{HSIC}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
-where $K$ is determined according to embedding semantic similarity and discussed in section B.3. Finally, the optimal bridge $L^{*}$ is selected by:
-$L^{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\text{-}L_{t})$ (5)
+Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy that measures the strongest dependency between a single neuron vector $\mathbf{x}_{i}\in\mathbf{X}$ ($\mathbf{y}_{j}\in\mathbf{Y}$) and the entire distribution of the other, where $\mathbf{x}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$, computed as:
+$\mathrm{H}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=\frac{1}{2}\Big(\max_{i}\,\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\,\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\Big)$ (3)
+We compute the dependency scores layer by layer and average them across the middle $K$ layers to estimate the selection probability of $L_{y}$:
+$p(L_{y}|L_{s}\rightarrow L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{H}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
+where $K$ is determined according to embedding semantic similarity and discussed in section C.3. Finally, the optimal bridge $L^{*}$ is selected by:
+$L_{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\rightarrow L_{t})$ (5)
 4 Experiment
@@
 Implementation.
-We evaluate BridgeX-ICL on $2$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
-The evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages.
-Therefore, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw), and high-to-low (e.g., En-He) pairs, both within and across families.
-The classification of high-, moderate-, and low-resource languages is based on their proportion in LLM’s training corpora, following previous work Cieri et al.
+We evaluate BridgeX-ICL on $4$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
+As the evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw) and high-to-low (e.g., En-He) pairs, both within and across language families.
+The classification of high-, moderate-, and low-resource languages is based on their proportion in LLMs’ training corpora, following previous work Cieri et al.
 (2016).
-Since the bridge language should be well supported by LLMs, we take 6 languages in Indo-European family as candidate bridges.
+Since the bridge language should be well supported by LLMs, we select $6$ languages in the Indo-European family as candidate bridges.
+We also conducted an exploratory experiment using bridges not in the Indo-European family discussed in section 5.1.
 Datasets.
-To evaluate the generalization of bridge selection beyond the BLI task, we also evaluate on an additional cross-lingual task: Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al.
+To evaluate the generalization of bridge selection beyond the BLI task, we further evaluate cross-lingual Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al.
 (2024).
-To evaluate low-resource languages, a key challenge lies in lacking evaluation benchmarks.
-Although the ground truth MUSE Conneau et al.
-(2017) provides $110$ bilingual dictionaries for BLI task, it does not cover the tested 15 language pairs.
-To solve this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English.
+To verify robustness, we additionally consider two cross-lingual tasks: Cross-Lingual Question Answering (CLQA) and Cross-Lingual Natural Language Inference (XNLI).
+Since CLQA and XNLI cover only six of our evaluated language pairs, their experimental results are reported in Appendix B.
+To evaluate low-resource languages, a key challenge lies in the lack of evaluation benchmarks.
+Although the ground-truth MUSE Conneau et al.
+(2017) provides $110$ bilingual dictionaries for the BLI task, it does not cover the $15$ language pairs we tested.
+To address this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English.
 For languages not in MUSE (e.g., Swahili), we extracted word pairs from wiktionary_bli Izbicki (2022) to build En-Sw.
-We checked all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems.
-The constructed BLI dictionaries are available at: https://anonymous.4open.science/r/BLI--0481/.
+We verified all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems  [^1] The constructed BLI dictionaries are available at: https://github.com/xuyuemei/BLI-..
 Metrics.
-For BLI task, we use Precision@N metric, which measures the accuracy of the model’s top $N$ candidate translations.
+For the BLI task, we use the Precision@N metric, which measures the accuracy of the model’s top-$N$ candidate translations.
 In this study, $N$ is set to $1$.
-For the MRC task, we use accuracy to measure whether the model selects the correct answer from multiple choices.
+For the MRC task, we use accuracy to evaluate whether the model selects the correct answer from multiple choices.
 LLMs.
-We conducted experiments on two open-source LLMs: LLaMA-3-8B Grattafiori et al.
+We conducted experiments on two open-source LLMs: LLaMA-3-8B Dubey et al.
 (2024) and Mistral-7B-Instruct-v0.3 Jiang et al.
@@
 Their training corpora cover 176 and 53 languages, respectively, which include all the experimental low-resource languages and allow us to explore the underlying linguistic mechanisms.
-[…]
+Baselines.
+Baselines are divided into zero-shot, few-shot, and zero-shot with bridge.
+Specifically, zero-shot is the basic prompt setup, and few-shot builds on the zero-shot prompt by adding 1, 2, 3, or 4 samples.
+For zero-shot with bridge approach, we compare BridgeX-ICL with 5 baselines described below.
+1) Phylogenetic Distance Source/Target (Ph.D Source or Ph.D Target): Select the bridge language closest to the source or target language according to language genealogy of Glottolog Trees Hammarström et al.
+(2023).
+2) English Bridge: Use English as the bridge language.
+3) Sharing Matters: Wang et al.(2024) used activation values to find shared neurons across languages.
+We select language with the most shared neurons as the bridge language.
+4) IoU: Use Intersection over Union (IoU) Tan et al.
+(2024), also known as Jaccard index, to measure linguistic distance.
+Given neuron sets $\mathcal{T}_{u},\mathcal{T}_{v}$ associated with language $L_{u}$, $L_{v}$, $\mathrm{IoU}(\mathcal{T}_{u},\mathcal{T}_{v})=|\mathcal{T}_{u}\cap\mathcal{T}_{v}|/|\mathcal{T}_{u}\cup\mathcal{T}_{v}|$.
+Language with the highest average IoU score to $L_{s}$ and $L_{t}$ is selected.
+5) LAPE_overlap: Use entropy-based LAPE Tang et al.
+(2024) to identify language-specific neurons.
+We then compute cosine similarity on overlap neurons between the bridge and source/target.
+The language with the highest average similarity is selected.
+(a) Linguistic spectrum in LLaMA 3
+(b) Language similarity from Glottolog Trees
+Figure 4: Comparison of linguistic spectrum calculated based on overlapping neurons in LLaMA 3 and language similarity derived from Glottolog Phylogenetic Trees, including $15$ languages from 77 families.
+Darker blue indicates a higher language similarity.
 4.2 Main Results
 4.2.1 LLMs’ Linguistic Spectrum Discussion
-(a) LLaMA 3’s Linguistic spectrum
-(b) Human language similarity from Glottolog
-Figure 4: Comparison of linguistic spectrum calculated based on overlap neurons in LLaMA 3 and human language similarity derived from Glottolog Phylogenetic Trees, including 15 languages from 7 families.
-Darker blue indicates a higher language similarity.
-This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlap neurons in LLaMA 3 and Mistral, as presented in Figure 4 (a) and Figure 8 in B.2, respectively.
-To evaluate how closely the linguistic spectrum learned by LLMs align with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al.
-(2023), which encode hierarchical relationships among 8,000+ human languages, to draw human language similarity in Figure 4(b).
-The detailed linguistic similarity computed based on Glottolog is in Appendix A.
-In Figure 4, color intensity represents degree of similarity between languages, with darker blue indicating stronger degree and the diagonal is self-similarity (1.0).
+This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlapping neurons in LLaMA 3 and Mistral, as presented in Figure 4(a) and Figure 8 in Appendix C.2, respectively.
+To evaluate how closely the linguistic spectrum learned by LLMs aligns with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al.
+(2023), which encode hierarchical relationships among 8,000+ human languages, to derive human language similarity in Figure 4(b).
+The detailed process for computing linguistic similarity based on Glottolog Trees is in Appendix A.
+In Figure 4, darker blue indicates stronger similarity, and the diagonal denotes self-similarity (1.0).
+Linguistic spectrum learned by LLMs is not fully aligned with human language phylogeny.
+We observe strong neural similarities within language families, highlighted by a red text box in Figure 4, which matches human linguistic taxonomy.
+For example, Arabic (Ar) and Hebrew (He), within the Afro-Asiatic family, exhibit a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161 similarity.
+A similar pattern appears with Indonesian (Id) and Tagalog (Tl), both from the Austronesian family.
+In addition, high-resource Indo-European languages, such as French-Italian (Fr-It) and Portuguese-Spanish (Pt-Es), show the highest overlap scores, with the darkest blue in the bottom-right corner of the heatmap.
+But this alignment breaks down in high-to-low resource language pairs, and some cross-family pairs display unexpected high similarity scores, likely reflecting training data distribution rather than intrinsic linguistic relationships.
+LLMs build their own distinct understanding of language relationships.
+The calculated linguistic spectra of LLaMA 3 and Mistral are similar but not identical.
+The two models may choose different bridges for the same language pair, as discussed later.
+We observe that Arabic has the strongest similarity (0.479) with French in the Romance family, rather than with Hebrew (0.470) from its own Afro-Asiatic family.
+This counterintuitive result is likely due to the linguistic relationships learned by LLMs being primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al.
+(2023).
 Table 1: Comparison of BLI task improvement on $15$ language pairs.
 The highest gains are marked with bold in few-shot and zero-shot with bridge methods.
-’-’ indicates the selected bridge is either the source or target language.
+‘-’ indicates the selected bridge is either the source or target language.
 LLaMA-3-8B
@@
 Few-shot Four-shot +12.50 +20.50 +3.20 +7.00 +7.70 +15.80 +0.80 +3.20 +14.30 +22.00 +10.60 -6.00 +26.10 -3.70 +7.40
-Human source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
-Human target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
+Ph.D Source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
+Ph.D Target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
 English Bridge +10.80 +12.60 +11.40 +4.80 +10.70 +17.50 +10.10 +3.30 +3.60 +9.30 +10.40 +2.50 - - -
-Sharing matter +9.50 +14.50 +8.40 +2.60 +6.40 +17.10 +6.10 +3.90 +2.60 +15.20 +7.70 +2.40 +12.30 -3.30 -1.70
+Sharing Matters +9.50 +14.50 +8.40 +2.60 +6.40 +17.10 +6.10 +3.90 +2.60 +15.20 +7.70 +2.40 +12.30 -3.30 -1.70
 IoU Score +10.80 +12.60 +11.40 +4.80 +10.70 +13.20 +5.30 +3.50 +2.30 +15.20 +7.40 +3.80 +9.70 -6.10 -1.80
@@
 Few-shot Four-shot +3.20 +1.20 -0.50 +2.60 -2.90 -6.60 0.00 +2.20 +10.70 +4.10 +2.50 +2.60 +1.20 -2.00 +2.50
-Human source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
-Human target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
+Ph.D Source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
+Ph.D Target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
 English Bridge +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 - - -
-Sharing matter +7.90 +8.60 +6.40 +1.20 +4.60 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 -0.10 +1.50 +0.90
+Sharing Matters +7.90 +8.60 +6.40 +1.20 +4.60 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 -0.10 +1.50 +0.90
 IoU Score +0.50 +5.30 +4.90 +2.20 +5.90 +1.10 +7.50 +1.60 +2.40 -2.80 -0.50 +1.00 -1.00 +1.50 +2.20
@@
 Ours +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +2.40 -3.30 +3.30 +1.00 -0.10 +1.50 +2.00
-Linguistic spectrum learned by LLMs are not fully aligned with human languages.
-We can observe a strong neural similarity within language families, marked using red text box in Figure 4, which matches human linguistic taxonomy.
-For example, within Afro-Asiatic family, both from Arabic (Ar) and Hebrew (He), have a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161.
-In addition, high-resource Indo-European languages, such as Fr-It and Pt-Es, show the highest overlap scores, with darkest blue in the bottom-right corner of the heatmap.
-But this alignment breaks down between high- and low-resource languages, like Arabic-French with similarity $0.479$.
-LLMs build their own distinct understanding of language relationships.
-The calculated linguistic spectrum of LLaMA 3 and Mistral are similar but not the same.
-The two models may choose different bridges for the same language pair, as discussed later.
-This is likely because their internal linguistic relationships are primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al.
+4.2.2 Cross-lingual Results Analysis
+Table 1 compares the performance of BridgeX-ICL against various baselines on the BLI task across $15$ language pairs.
+We observe the following findings: 1) LLMs exhibit poor and imbalanced performance on low-resource languages.
+For example, LLaMA 3 achieves its best BLI performance of $69.90$ on the Ar-Ja pair, but its worst of $25.90$ on the Id-Sw pair.
+2) LLMs are few-shot multilingual learners, as also noted in Al Nazi et al.
 (2024).
-This explains why Arabic has the strongest similarity ($0.479$) with French in Romance family, rather than with Hebrew ($0.470$) in the same Afro-Asiatic.
-4.2.2 Cross-lingual Results Analysis
-This section compares performance of BridgeX-ICL against various baselines on BLI task across $15$ language pairs, as presented in Table 1.
-Table 2: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selection and bold marks the highest gains at each language pair.
+However, few-shot X-ICL does not consistently yield stable gains.
+For example, one-shot sometimes performs worse than zero-shot, and performance often stabilizes or may decline once the number of shots exceeds $3$.
+This suggests that when applying few-shot X-ICL, 3-shots will be enough.
+3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages.
+BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method.
+While phylogenetic distance source/target methods are the least effective.
+It seems using English as the default bridge is cost-effective, which will be discussed in section 5.2.
+5 Discussion
+5.1 Candidate Bridge Language Selection
+To validate the rationale for selecting Indo-European languages as bridge candidates, Table 2 presents an exploratory experiment in which all languages in our study were evaluated as potential bridges, using $6$ representative language pairs.
+It shows that Indo-European languages on average outperform the nine non-Indo-European languages.
+Interestingly, some non-Latin-script languages, such as Chinese, also demonstrate potential as effective bridges.
+The results provided valuable insights: only languages well supported by LLMs functioned as effective bridges.
+Based on this observation, we selected $6$ languages in Indo-European as the final candidates in our study.
+Table 2: Candidate bridge language selection comparing Indo-European languages with 99 non-Indo-European languages. ’-’ indicates the selected bridge is either the source or target language.
+Zero-shot Zh Ja Ar He Id Tl Fi Hu Sw Indo-European (Avg.)
+Zh-Ja 67.10 - - +3.30 +0.80 0 -5.10 +2.80 +5.20 -9.90 +9.43
+Zh-He 44.10 - +7.90 +9.60 - +4.80 -0.20 +6.70 +10.80 -8.80 +14.37
+Ar-Tl 46.70 +4.50 +3.10 - -7.40 -7.10 - +3.40 +4.90 -6.10 +6.88
+Ar-He 47.00 +14.90 +11.70 - - -5.20 +1.40 +11.60 +15.30 -3.00 +15.72
+Id-Ja 62.50 -2.40 - -9.00 -5.20 - -11.90 -12.10 -9.60 -24.80 +1.40
+Id-Sw 25.90 +3.20 +1.50 +1.20 -0.80 - -3.90 -0.50 +1.30 - +3.43
+5.2 Application of Bridge Language
+Beyond the BLI task, Table 3 evaluates BridgeX-ICL on the MRC cross-lingual task.
+The prompt for zero-shot with bridge in MRC is detailed in Appendix F.
+Appendix B further evaluates BridgeX-ICL on CLQA and XNLI cross-lingual tasks.
+Results show our approach works well and benefits more from LLaMA 3 than from Mistral.
+For example, BridgeX-ICL improves the performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across $15$ language pairs, while the average improvement on Mistral is 4.484.48%.
+English is selected as the optimal bridge in $9$ out of $15$ language pairs in LLaMA 3 ($6$ out of $15$ in Mistral).
+This is partly due to the unbalanced language abilities of LLMs across 5 candidate Indo-European bridges.
+As discussed in Figure 3, another key factor is the model’s inherent preference for English-pivot during cross-lingual transfer.
+Table 3: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selections, and bold marks the highest gains at each language pair.
 LLaMA-3-8B Mistral-7B
@@
 En-Sw Fr 64.60 +7.40 Fr 51.20 +0.20
-We can find: 1) LLMs exhibit poor and imbalanced performance on low-resource languages.
-For example, LLaMA 3 achieves its best BLI performance of $69.90$ on Ar-Ja pair, but worst of $25.90$ on Id-Sw pair.
-2) LLMs are few-shot multilingual learners.
-However, few-shot X-ICL does not consistently yield stable gains, e.g., one-shot falling below zero-shot, and will not lead to further gains when the number of shots goes beyond $3$.
-It indicates when applying few-shot X-ICL, 3-shots will be enough.
-3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages.
-BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method.
-While human source/target methods are the least effective.
-It seems using English as the default bridge is cost-effective, which will be discussed in section 5.1.
-5 Discussion
-5.1 Application of Bridge Language
-Beyond BLI task, Table 2 evaluates BridgeX-ICL on the MRC cross-lingual task.
-The prompt for zero-shot with bridge in MRC is detailed in Appendix E.
-Results show our approach still works well on MRC task and benefits more on LLaMA 3 than Mistral.
-For example, BridgeX-ICL improves performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across 15 language pairs, while the average improvement is 4.484.48% on Mistral.
-English is selected as the optimal bridge in 9 out of 15 language pairs.
-The reason is partly due to LLMs’ unbalanced language abilities across 5 candidate bridges.
-As discussed in Figure 3 another key factor is LLMs’s inherent preference of English-pivot during cross-lingual transfer.
 5.3 Ablation Study
-In this part, we conduct ablation study to evaluate the impact of neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task for example.
-Table 3: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from FLORES+ dataset.
+In this part, we conduct ablation study to evaluate the impact of the constructed neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task as an example.
+Table 4: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from the FLORES+ dataset.
 LLaMA-3-8B Mistral-7B
@@
 En-Sw 26.20  26.20   10.20  10.30  $\uparrow$
-Table 3 presents the results of ablation experiments to compare “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al.
+Table 4 presents the ablation results by comparing “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al.
 (2024).
-Appendix C illustrates the detailed construction for “w/o *”.
-These results highlight impact of probing data in neuron manipulation.
-Table 4 in Appdedix C compares the performance of HSIC and Cosine similarity and shows the proposed HSIC helps to capture the dependency between language-overlapping neurons and specific neurons.
+The detailed construction of “w/o *” is provided in Appendix D.
+The results highlight the crucial impact of probe data on effective neuron manipulation.
+Furthermore, Table 6 in Appendix D compares HSIC with Cosine similarity, showing that HSIC better captures the dependency between language-overlapping neurons and specific neurons.
 5.4 Overlapping Neuron Distribution
@@
 Figure 5 compares neurons in language pairs within the same family (e.g., Ar-He) and across families (e.g., Ar-Sw).
-Obviously, Ar-He shares more overlap neurons.
-Similar observations can be found in comparing language pairs from different source languages (Zh, Ar, Id, and En ) to a same target language He (Figure 10 in B.4) From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages.
+Obviously, Ar-He shares more overlapping neurons.
+Similar observations can be found in comparing language pairs from different source languages to a same target language (Figure 10 in Appendix C.4)
+From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages.
 Obviously, low-resource languages within the Uralic family have the highest activation frequency, while Indo-European languages have the lowest.
@@
 We propose a simple yet effective language-bridge approach with the help of neuron interpretation.
-To ensure accurately and fully activate overlap neurons across languages, we construct neurons probe data from the ground-truth MUSE dictionaries.
-By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate the efficacy and generalization of our approach.
+To ensure accurate and full activation of overlap neurons across languages, we construct probing data from the ground-truth MUSE dictionaries.
+By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate its efficacy and generalization.
 Limitations
-This work focuses on sharing neurons across languages and relies on evaluated data to validate the effectiveness of our approach.
-Due to the limitation of lacking evaluation benchmarks for low-resource languages, the experiments are conducted on 15 language pairs and have not been extensively validated on a wide range of low-resource languages.
-We select 4 typologically diverse low-resource languages from distinct families and test their performance on 2 cross-lingual tasks.
-Second, our study reveals that high-quality probe data is essential for accurately analyzing neurons’ behaviors of low-resource languages.
-The proposed linguistic distance measurement is probe-data-induced, providing qualitative insights, but lacks quantitative precision.
-Acknowledgments
-This work was supported by the National Social Science Foundation (No.24CYY107) and the Fundamental Research Funds for the Central Universities (No.2024TD001).
+This work focuses on sharing neurons across languages and relies on the evaluated datasets to validate the effectiveness of our approach.
+Due to the lack of comprehensive benchmarks for low-resource languages, our experiments cover only $15$ language pairs and select $4$ low-resource languages from distinct families as target languages and test their performance on $4$ cross-lingual tasks.
+Second, our study reveals that high-quality probe data is crucial to accurately analyze neuron behaviors of low-resource languages, while the proposed linguistic distance measurement is probe-data-induced, offering qualitative but not quantitative insights.
+Finally, although bridge selection should ideally follow linguistic phylogeny, we aim to select bridges that LLMs can best exploit, inevitably reflecting their training biases.
 Appendix A Appendix: Linguistic Similarity Based on Glottolog Phylogenetic Trees
-We leverage Glottolog verion 5.1 Hammarström et al.
+We leverage Glottolog version 5.1 Hammarström et al.
 (2023) as a foundational phylogenetic framework to calculate the linguistic similarity of human languages.
@@
 Data Preprocessing.
-The preprocessing pipeline consists of three steps: 1) Glottocode Identifiers with regex pattern matching to ensure unambiguous language node identification.
-2) Standardize node naming with underscores (e.g., [sini1245]→ _sini1245_), ensuring consistent formatting in downstream phylogenetic analyses.
+The preprocessing pipeline consists of three key steps: 1) Locate Glottocode identifiers with regex pattern matching to ensure unambiguous language node identification.
+2) Standardize node naming with underscores (e.g., [sini1245] $\rightarrow$ _slaini1245_), ensuring consistent formatting in downstream phylogenetic analyses.
 3) Mitigate encoding conflicts through temporary file caching.
-These steps preserve both accurate phylogenetic tree parsing and computational compatibility.
-Similarity Calculation.
-The proposed metric integrates two well-established phylogenetic principles from historical linguistics, which includes node distance normalization and depth-adjusted compensation.
+These steps preserve accurate parsing of phylogenetic tree while maintaining computational compatibility.
+Similarity Calculation. The proposed metric integrates two well-established principles from historical linguistics: node distance normalization and depth-adjusted compensation.
 First, building upon Wichmann & Holman’s framework for typological stability assessment Wichmann and Holman (2009), we compute the inter-language distance $d(L_{1},L_{2})$ between languages $L_{1}$ and $L_{2}$ using ETE3’s optimized tree traversal algorithms Huerta-Cepas et al.
-(2016) and then normalize it to make distances comparable across language families, calculated as:
+(2016).
+We then normalize the distance to make it comparable across language families, calculated as:
 […]
@@
 $\mathrm{Sim}(L_{1},L_{2})=S_{\mathrm{distance}}\times\alpha_{\mathrm{depth}}$ (8)
-Appendix B Appendix: Neuron Patterns
-B.1 Deactivation Overlap Neurons
-Figure 7 presents overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
-Figure 7: Overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
-B.2 Linguistic Spectrum in Mistral
-Figure 8: Mistral’s linguistic spectrum across 15 languages from 7 families. The color intensity represents the degree of overlap between language pairs.
-B.3 Parameter $K$ Discussion
+Appendix B Appendix: Supplementary Cross-lingual Tasks
+To further verify robustness, we conduct additional experiments using the bridge languages selected by our method on downstream tasks, including Cross-Lingual Question Answering (CLQA) and Cross-Lingual Natural Language Inference (XNLI), as shown in Table 5.
+Due to the limited availability of cross-lingual benchmarks covering our target low-resource languages (e.g., Tagalog), the evaluation is restricted to $6$ language pairs for CLQA and $3$ pairs for XNLI.
+Table 5: Evaluation on CLQA and XNLI cross-lingual tasks. Bold highlights improved performance.
+CLQA XNLI
+Zero-shot Ours Zero-shot Ours
+Zh-Ja 42.80 -0.60 - -
+Zh-Sw 38.20 +5.20 35.90 +1.30
+Ar-Ja 41.60 +6.20 - -
+Ar-Sw 34.20 +2.40 35.50 +3.70
+Id-Ja 34.60 +3.60 - -
+Id-Sw 34.80 +5.20 36.60 +0.60
+Appendix C Appendix: Neuron Patterns
+C.1 Deactivation Overlap Neurons
+Figure 7 presents the distribution of overlap neurons and their deactivation effects on the Chinese-Hebrew (Zh-He) BLI task.
+Figure 7: Overlap neuron distributions and their deactivation effects on the Chinese-Hebrew BLI task.
+C.2 Linguistic Spectrum in Mistral
+Figure 8: Mistral’s linguistic spectrum across $15$ languages from 77 families. The color intensity represents the degree of overlap between language pairs.
+C.3 Parameter $K$ Discussion
 Here we discuss which $k$ middle layers should be selected to quantify linguistic similarity.
-According to observations in section 3.3.3, neurons in middle-layers should be prioritized over final-layers when measuring language similarity.
-We use embedding semantic similarity metric to determine $K$.
-We compute the layer-wise embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew pairs when predicting the same Hebrew token.
-As presented in Figure 9, we find embedding similarity is stable in the middle layers 10-21 and is not sensitive to variations in the predicted tokens.
+According to observations in section 3.3.3, neurons in middle layers should be prioritized over final layers when measuring language similarity.
+We use the embedding semantic similarity metric to determine $K$.
+For example, we analyze Arabic-Hebrew and Chinese-Hebrew translation pairs by prompting LLMs with the same semantical words in Arabic and Chinese to generate the corresponding Hebrew translation.
+We then compute the layer-wise embedding semantic similarity between the two pairs and identify layers in which this similarity is insensitive to variations in the predicted tokens, reflecting the inherent distance between the languages.
+As shown in Figure 9, we find embedding similarity is stable in the middle layers 10-21 of LLaMA 3 and is not affected by token-level variations.
 Therefore, $K$ layers is set to be 10-21 in LLaMA 3 and 15-23 in Mistral.
-Figure 9: Embedding semantic similarity between Arabic-to-Hebrew and Chinese-to-Hebrew translations when predicting the same token at each layer.
-B.4 Overlap Neuron Distribution
+Figure 9: Embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew translations when predicting the same token at each layer of LLaMA 3.
+C.4 Overlap Neuron Distribution
 Figure 10 presents the distribution of overlap neurons across different language pairs in LLaMA 3.
 Figure 10: Distribution of overlap neurons across different language pairs in LLaMA 3.
-Appendix C Appendix: Ablation Results
-This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on bridge selection.
+Appendix D Appendix: Ablation Results
+This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on the bridge language selection.
 Table 4 and Table 6 present the results of ablation experiments on the BLI task.
 “w/o *” denotes replacing our probe data with a simplified version based on FLORES+.
-For example, “w/o *” probe data in Indonesian-Hebrew is illustrated in Figure 11.
+For example, Figure 11 illustrates the construction of “w/o *” probe data for Indonesian-Hebrew.
 Table 6: Performance comparison of using HSIC and Cosine similarity metrics on the BLI task.
@@
 En-Sw Fr 26.20 $\uparrow$ It 21.50
-Figure 11: Example of “w/o *” probe data in Indonesian-Tagalog.
-Appendix D Appendix: Neuron Semantic Analysis
-In this section, we analyze the semantic similarity of overlapping neurons in two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, different language family) and Portuguese-Spanish-Italian (pt-Es-It, same language families).
-For each group, we select overlapping neurons and the same number of randomly sampled neurons for comparison.
-We input $m$ parallel sentences and record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for overlap and random neurons.
+Figure 11: Example of “w/o *” probe data in Indonesian-Hebrew.
+Appendix E Appendix: Neuron Semantic Analysis
+In this section, we analyze the semantic similarity of overlapping neurons across two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, from different language families) and Portuguese-Spanish-Italian (Pt-Es-It, from the same language family).
+Our goal is to examine whether overlapping neurons cluster together within the same language family or also across different families.
+For comparison, we select $100$ overlapping neurons and randomly sample the same number of non-overlapping neurons for comparison.
+We feed the model with $m$ parallel sentences and then record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for both overlapping and random neurons.
 These matrices are mapped to a 2D semantic space using UMAP McInnes et al.
-(2018), with each point representing a neuron activated by a sentence.
-As shown in Figure 12, randomly sampled neurons align with linguistic relationships.
-For example, random neurons of He and Tl are close, indicating semantic proximity in the same family.
-The overlap neurons cluster together, both within and across language families, proving that our approach can effectively capture semantic similarity.
+(2018), with each point representing a neuron activated by $m$ sentences from a language.
+Colored circles of red, yellow, and blue denote overlapping neurons and colored $\times$ symbols of red, yellow, and blue denote random neurons.
+As presented in Figure 12, the overlapping neurons, whether identified within the same family or across language families, cluster closely together.
+This suggests that our approach can effectively identify neurons that encode genuine shared semantics across languages.
+In addition, random neurons tend to align with linguistic relationships, clustering when feeding LLMs tokens from topologically related languages, such as Portuguese (Pt), Spanish (Es), and Italian (It), while remaining dispersed when feeding LLMs tokens from distant languages, like Hebrew (He), Tagalog (Tl), and Swahili (Sw).
+(a) Neuron semantic similarity in Pt-Es-It
+(b) Neuron semantic similarity in He-Tl-Sw
 Figure 12: Visualization of semantic similarity comparing language-overlapping neurons and randomly sampled neurons.
-Appendix E Appendix: Prompt Templates
+Appendix F Appendix: Prompt Templates